40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jan 4 02:58:53 2011, Jenne, Update, IOO, Power into Mode Cleaner increased 
    Reply  Tue Jan 4 11:06:32 2011, Koji, Update, IOO, Power into Mode Cleaner increased 
       Reply  Tue Jan 4 13:57:48 2011, kiwamu, Update, IOO, incident mirror changed from 45deg to 0 deg 
          Reply  Wed Jan 5 03:06:02 2011, kiwamu, Update, IOO, beam spots on MC mirrors 
       Reply  Tue Jan 4 18:37:18 2011, Jenne, Update, IOO, MCWFS aligned 
Message ID: 4105     Entry time: Tue Jan 4 13:57:48 2011     In reply to: 4104     Reply to this: 4110
Author: kiwamu 
Type: Update 
Category: IOO 
Subject: incident mirror changed from 45deg to 0 deg 

I replaced the final steering mirror (Y1-1037-45-S) in the zig-zag path on the PSL table by a 0 deg mirror Y1-1037-0.

With a sensor card I confirmed the transmission reduced a lot after the replacement.

As we expected, the replacement of the mirror caused a mis-alignment of the incident beam axis to the MC, so I compensated it by touching the angle of the mirror a little bit.

After the alignment of the mirror, the MC is still resonating at TEM00.

We will check the spot positions more accurately by A2L technique.

Quote:

- Y1-45S has 4% of transmission. Definitively we like to use Y1-0 or anything else. There must be the replaced mirror. 

I think Suresh replaced it. So he must remember wher it is.

ELOG V3.1.3-